OpenStudy (anonymous):

write the equation in slope intercept form of the line that has a slope of 2 and contains the point (3,7)

OpenStudy (lgbasallote):

\[y - y_1 = m(x-x_1)\] \[y - 7 = 2 (x-3)\] \[y - 7 = 2x - 6\] \[y = 2x - 6 + 7\] \[y = 2x + 1\]

OpenStudy (anonymous):

thank you